OpenStudy (anonymous):

d to the fourth power when d = 1/3 can someone help me solve this?

OpenStudy (dumbcow):

(1/3)^4 = 1^4/3^4 = 1/(3^4)

OpenStudy (radar):

Is d the number fraction 1/3? If so you raise the numerator and the denominator to the fourth power.

OpenStudy (anonymous):

so 1/81 is not correct

OpenStudy (dumbcow):

no that would be correct

OpenStudy (radar):

I believe that should be your final answer, 1 to the 4th power is 1 and 3 to the fourth power is 81

OpenStudy (anonymous):

ok cool than i'm on the right track Thanks

OpenStudy (anonymous):

which is the base? 1/3?

OpenStudy (radar):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (radar):

NP & GL